# Atom-Induced Field Squeezing Predicted by Magnus Expanding the Jaynes-Cummings Model for a Two-Level Atom

### v2

@@ IV Conclusion
In this paper, we kept all the terms of the Jaynes-Cummings Hamiltonian by not performing the RWA prematurely. We performed first and second order Magnus Expansion for the time-evolution operator using the Jaynes-Cummings Hamiltonian rotated in both atom and field frames. We have seen why the RWA is valid from the first-order Magnus expansion. On the other hand, we have seen the non-trivial contributions from the counterrotating terms, such as the Bloch-Siegert Shift and the creation of squeezed light due to the presence of atoms. The key contribution of this paper is the analytic prediction of field squeezing from a two-level atom in the JC model using the Magnus expansion. This term, emerging beyond the RWA, offers a new theoretical insight into cavity QED systems.
@@
Our long-term goal is to physically explain the nature of RWA by finding its deviations from the exact Magnus Expansion approach. This work highlights how the Magnus Expansion not only justifies the use of the RWA but also reveals non-negligible quantum effects, offering a physical framework to understand the effects beyond the RWA.
Acknowledgements.
I would like to express my appreciation to Professor Daniel F. V. James from the Department of Physics, University of Toronto, for reading the manuscript and providing helpful discussions and comments.
Disclosure of Interest
The author reports that there are no competing interests to declare.
Appendix A BCH identities
For arbitrary operators $\hat{A},\hat{B}$ and an arbitrary complex number $\zeta$, the Baker-Campbell-Hausdorff (BCH) lemma states the following identity.
$exp(\zeta\hat{A})\ \hat{B}\ exp(-\zeta\hat{A})=\hat{B}+\zeta[\hat{A},\hat{B}]+\frac{\zeta^{2}}{2!}[\hat{A},[\hat{A},\hat{B}$]]+...
Using the BCH lemma, we derive the following identities.
$\hat{V}^{\dagger}\hat{a}^{\dagger}\hat{V}$ $=exp(i\omega\hat{a}^{\dagger}\hat{a}t)\hat{a}^{\dagger}exp(-i\omega\hat{a}^{\dagger}\hat{a}t)$
$=\hat{a}^{\dagger}+i\omega t[\hat{a}^{\dagger}\hat{a},\hat{a}^{\dagger}]+\frac{(i\omega t)^{2}}{2}[\hat{a}^{\dagger}\hat{a},[\hat{a}^{\dagger}\hat{a},\hat{a}^{\dagger}]]$
Since $\hat{a}^{\dagger}\hat{a},\hat{a}^{\dagger}]=\hat{a}^{\dagger}$, we get the following series.
@@ Appendix A BCH identities
Doing the same for the annihilation operator, we get the following identity.
[…]
Using the Taylor series expressions for exponents, we get the necessary expressions to perform the light-rotations as follows.
$\hat{V}^{\dagger}\hat{a}^{\dagger}\hat{V}=\hat{a}^{\dagger}e^{i\omega t}$ (32)
$\hat{V}^{\dagger}\hat{a}\hat{V}=\hat{a}e^{-i\omega t}$ (33)
Appendix B Proof that the Coefficient of the Squeezing Term does not Diverge on Resonance
Recall that the coefficient of the squeezing term is as follows.
$\zeta=\frac{g^{2}}{2}\frac{\omega_{0}e^{2i\omega t}-\omega e^{i\Sigma t}+\omega e^{i\Delta t}-\omega_{0}}{(\omega^{2}-\omega_{0}^{2})\omega}$ (34)
On resonance, $\omega=\omega_{0}$. This means $\Sigma=\omega+\omega_{0}=2\omega$ and $\Delta=\omega-\omega_{0}=0$. On the limit of resonance, the constant can be represented as follows.
$\zeta$ $=\frac{g^{2}}{2}\frac{(\omega_{0}-\omega)e^{2i\omega t}+(\omega-\omega_{0})}{\omega(\omega-\omega_{0})(\omega+\omega_{0})}$
$=\frac{g^{2}}{2}\frac{(\omega-\omega_{0})(1-e^{2i\omega t})}{\omega(\omega-\omega_{0})(\omega+\omega_{0})}$ (35)
On a limit that $\omega$ approaches $\omega_{0}$, the terms $\omega-\omega_{0}$ cancel and the constant reduces to a non-divergent form on resonance.
$\lim_{\omega\to\omega_{0}}\zeta=\frac{g^{2}}{2}\frac{1-e^{2i\omega t}}{\omega(\omega+\omega_{0})}$ (36)
Since $\zeta$ does not diverge on resonance, its complex conjugate $\zeta^{*}$ does not diverge either.

### v4

@@ IV Conclusion
In this work, we kept all the terms of the Jaynes-Cummings Hamiltonian by not performing the RWA prematurely. We performed first and second order Magnus Expansion for the time-evolution operator using the Jaynes-Cummings Hamiltonian rotated in both atom and field frames. We have seen the non-trivial contributions from the counterrotating terms, such as the Bloch-Siegert Shift and the creation of squeezed light due to the presence of a two-level atom. The key contribution of this paper is the analytic prediction of field squeezing from a two-level atom in the JC model using the Magnus expansion. Such terms, emerging beyond the RWA, not only offers a new theoretical insight into engineering squeezed states of light but also opens doors for QND measurements using simple systems. Moreover, we see that the energy shifts such as the AC-Stark shift, Bloch-Siegert shift, and the vacuum energy shift are time-dependent, by a function of f(t). It is interesting to go beyond the second order to see what other non-trivial contributions can be witnessed by the non-RWA term. It is also interesting to extend this research and determine up to second-order Magnus expansion using general multilevel atomic systems under the dipole selection rules. Future research should also study the feasibility of using a two-level system for QND measurements.
The prediction of a squeezing coefficient directly from the JC Hamiltonian emphasizes the broader significance of developing theoretical quantum optics from first principles. By deriving squeezing behavior without recourse to phenomenological assumptions, this work reinforces the fundamental link between atom–field interactions and nonclassical light. While a single two-level atom cannot serve as a practical source of squeezed states, the analytical framework presented provides a clear theoretical quantification of the atom-induced squeezing coefficient. Beyond its immediate implications, this approach highlights that rigorous first-principles treatments not only sharpen our understanding of quantum optical processes but also connect to the well-studied literature on the SU(1,1) representation of field squeezing and fluctuations.
Appendix A Commutation Relations of Two-Photon Operators
One normally expects the higher-order terms in a perturbative series to have small corrections to the squeezing coefficient. However, the non-commuting structure embedded in the second order itself induces corrections to the squeezing coefficient without needing to go to higher orders. To demonstrate this, recall the Baker-Campbell-Hausdorff(BCH) theorem [26].
@@ Appendix A Commutation Relations of Two-Photon Operators
Notice that the commutation relations of the individual field operators in equation 25 are as follows.
$[\frac{1}{2}(\hat{a}^{\dagger}\hat{a}+\frac{1}{2}),\frac{1}{2}\hat{a}^{2}]=-\frac{1}{2}\hat{a}^{2}$ (28)
$[\frac{1}{2}(\hat{a}^{\dagger}\hat{a}+\frac{1}{2}),\frac{1}{2}(\hat{a}^{\dagger})^{2}]=\frac{1}{2}(\hat{a}^{\dagger})^{2}$ (29)
$[\frac{1}{2}(\hat{a}^{\dagger})^{2},\frac{1}{2}\hat{a}^{2}]=-(\hat{a}^{\dagger}\hat{a}+\frac{1}{2})$ (30)
When expanding the exponent using the BCH theorem, one quickly starts noticing that every commutator one evaluates, extra terms of $(\hat{a}^{\dagger})^{2},\hat{a}^{2},\hat{a}^{\dagger}\hat{a}$ are produced. The result is a never-ending loop with every term providing extra squeezing coefficients. However, there is a compact way to evaluate the exponent since the field operators form an SU(1,1) group.
Three operators $\{\hat{K}_{0},\hat{K}_{\pm}\}$ forms an SU(1,1) group if obey the following commutation relations [17].
$[\hat{K}_{0},\hat{K}_{\pm}]=\pm\hat{K}_{\pm}$ (31)
$[\hat{K}_{+},\hat{K}_{-}]=-2\hat{K}_{0}$ (32)
Defining $\hat{K_{0}}=\frac{1}{2}(\hat{a}^{\dagger}\hat{a}+\frac{1}{2})$, $\hat{K}_{+}=\frac{1}{2}(\hat{a}^{\dagger})^{2}$ and $\hat{K}_{-}=\frac{1}{2}\hat{a}^{2}$, and we can see that the operators form an SU(1,1) group. Then we can rewrite equation 25 as follows by defining $\alpha(t)=2ig^{2}f(t)\hat{\sigma}_{z}$, and $\beta(t)=-\hat{\sigma}_{z}\zeta(t)$.
@@
It is shown in Truax’s paper that equation 33 can be disentangled into the following form[26].
$e^{\hat{\Omega}_{2}(t)}=$ $\exp\left[p_{2}\hat{K}_{+}\right]\exp[p_{0}\hat{K}_{0}]\exp[p_{1}\hat{K}_{-}$], (34)
The coefficients $p_{i}(\alpha(t),\beta(t))$ were evaluated in explicit closed forms in Truax’s article [26]. Following Truax’s discussion, we conclude that the second-order Magnus operator resolves in a compact, closed-form unitary evolution.